The concentration of hydroxyl ion in solution left after mixing 100 mL of 0.1 M MgCland 100 mL of 0.2 M NaOH [ Ksp of Mg(OH)2 = 1.2 × 10-11]  is
​a.  2.8 × 10-3 M
​b.  2.8 ×10-2 M
​c.  2.8 ×10-4 M
​d.  2.8 ×10-5 M
Ans ;- (C)
Please explain in detail.........

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii 
The reaction can be written as
                        MgCl2 + 2NaOH =.>  Mg(OH)2 + NaCl
mM Before           10           20                0                0
mM After                0            0                  10             20
 
The product of [Mg2+][OH-] is therefore
10/100 * (20/200)^2 = 5* 10^-4
Solublity of Mg(OH)2 is derived by 
S= (Ksp/4)1/3 
S= 1.4 * 10^-4
[OH-] = 2S = 2.8 * 10^-4 …...............(c)